You are planning to do a PCR with a DNA sample that you got an OD260 of 0.02 Absorbance after doing a 1ul to 99ul dilution.  Using the following formula, calculate the concentration of DNA: OD260 * 50ug/ml * dilution factor.  (Be careful of units).

A.

1 ug/ul

B.

0.1ug/ul

C.

10 ug/ul

D.

100 ug/ul
Answer #1

We know that the

Dilution Factor = (Total Volume of Sample)/(Volume of sample)

Dilution Factor = [(VDNA sample)+(VBuffer)]/(V​​​​​​DNA sample)

Given, VDNA sample = 1 ul

VBuffer = 99 ul

Dilution factor = [(1 ul)+(99 ul)]/(1 ul) = (100 ul)/(1ul) = 100

Given,

OD260 = 0.02

Putting these values in the given formula, we get

The concentration of DNA = OD260 * 50ug/ml * dilution factor = 0.02 * 50 ug/ml * 100 = 100 ug/ml = 100 ug/(1000 ul) = 0.1 ug/ul

[ 1 ml = 1000 ul]

Hence, answer to this question is (B) 0.1 ug/ul